# Overcoming Joint Intractability with Lossless Hierarchical Speculative Decoding

### v1

@@ 1 Introduction
Inference speed has become paramount for Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Bai et al., 2023), which generate text auto-regressively. Recent advances in test-time scaling (OpenAI, 2024; Guo et al., 2025) have further underscored its importance. While techniques like pruning (Frankle and Carbin, 2018; Sun et al., 2023a) and quantization (Shen et al., 2020; Xiao et al., 2023) improve efficiency but sacrifice performance, Speculative Decoding (Leviathan et al., 2023) achieves speedups while preserving the target model’s distribution, making it a particularly appealing alternative. It adopts a smaller model to make proposals and a larger model to select from them with a grounded verification strategy. Most approaches prioritize the drafting phase, but further gains face diminishing returns. Driven by the verification bottleneck, recent methods (Cai et al., 2024; Zhou et al., 2024; Narasimhan et al., 2024) trade off fidelity for speed, relying on task-specific tuning; their performance typically remains constrained to carefully curated scenarios.
@@ 2 Related Work
Drafting Phase. Drafting methods can be grouped into three categories: (1) Single-draft. Early SD methods (Leviathan et al., 2023) inspired PaSS (Monea et al., 2023) and Draft&Verify (Zhang et al., 2024), improving efficiency via multi-token generation or selective layer skipping. GLIDE (Du et al., 2024) (shared KV-cache) offers further speedups but requires task-specific tuning. (2) Retrieval-based. LLM-A (Yang et al., 2023) and ReST (He et al., 2023) generate drafts from reference texts, potentially reducing latency, but face database limitations, distribution gaps, and reliance on greedy decoding. (3) Multi-draft. Tree-attention frameworks—SpecInfer (Miao et al., 2024), Medusa (Cai et al., 2024), and Eagle (Li et al., 2024)—expand many branches, quickly exhausting memory. Medusa and Eagle also predict drafts from the target model’s hidden features rather than a separate draft model, further boosting speed but requires task-specific tuning.
Verification Phase. Verification methods trade fidelity for speed. Lossless approaches (Sun et al., 2023b; Yang et al., 2024; Hu et al., 2025) guarantee exact recovery but are costly. Block Verification (Sun et al., 2024) partially alleviates this bottleneck but offers limited improvement and low interpretability and integrity. Lossy methods—including BiLD (Kim et al., 2023), MTAD (Qin et al., 2025), DistillSpec (Zhou et al., 2024), Medusa-2 (Cai et al., 2024), and SpecCascade (Narasimhan et al., 2024) increase speed but compromise distribution fidelity and require task-specific tuning. In addition, Medusa and EAGLE always accept the first draft token to improve throughput, trading off exact recovery of the target distribution.
@@ 5.3 Computational Efficiency
The verification stage in HSD adds negligible overhead compared to the savings from reduced target model forward passes. Thanks to parallelized computations across both the vocabulary and draft positions, HSD is nearly as efficient as tokenwise verification. Runtime measurements (Appendix G) show that verification accounts for less than 1% of total decoding time, with the majority still spent on draft and target forward passes. These results demonstrate that HSD is not only theoretically lossless but also practically efficient, as further confirmed by our experiments in Section˜6.
@@ 6.1 Experiment Setting
Baselines and Metrics. We compare two lossless verification methods—Token-wise (Leviathan et al., 2023) and Block-wise (Sun et al., 2024)—using two metrics: Block Efficiency (tokens/step) and Decoding Speed (tokens/second). Block Efficiency measures the average tokens generated per serial call to the target model, reflecting intrinsic efficiency independent of hardware. Decoding Speed indicates tokens produced per second for practical reference, though it depends on hardware and implementation. Additional details and extended evaluations are in Section˜E.
@@ 6.2 Experiment Results
Main results.  Table 1 summarizes the performance of HSD across datasets and model scales using the Qwen2.5 suite (0.5B as draft,14B, 32B, and 72B as targets). Overall, HSD consistently improves both Block Efficiency (BE) and Decoding Speed (DS) relative to Tokenwise and Blockwise verification. For GSM8K, the gains are stable across scales, with BE improvements of 5.2%–5.4% at 14B/32B and 3.3% at 72B, accompanied by DS increases of up to 10.7%. On HumanEval, the effect is more pronounced: BE rises by 9.5% and 12.3% at 14B and 32B, while DS improves by 9.3% and 11.4%; even at 72B, HSD maintains positive margins (3.3% BE, 4.5% DS). For CNN/DailyMail, the improvements are moderate but consistent, with BE gains of 4.2%–8.4% and DS gains of 3.4%–7.2%. Taken together, these results demonstrate that HSD not only outperforms Tokenwise verification but also provides consistent advantages over Blockwise verification, yielding average improvements of approximately 6.2% in BE and 6.7% in DS. The consistency of these gains across datasets and scales highlights the robustness and scalability of the approach.
Multi-draft. To demonstrate the advantage of HSD, we compare it with token-wise speculative decoding in a multi-draft setting. For simplicity—and without loss of generality—we adopt Recursive Reject Sampling (RRS) with replacement (Yang et al., 2024) as the baseline for its scalability and independence from complex tree attention mechanisms. Notably, since it is not straightforward to extend blockwise verification to the multi-draft setup, we omit it from our comparison. We evaluated multi-draft generation with 11 candidate drafts in Table 2, and HSD yields an average 5.9% improvement in Block Efficiency and 4.7% improvement in Decoding Speed over token-wise decoding. These results further underscore the strong potential of HSD to improve performance when combined with complementary or orthogonal techniques.
@@
(a) Comparison of different sampling temperatures. The draft length $\gamma$ is set to 10.
@@
(b) Comparison of different draft lengths. The temperature is set to 1.
@@
Extended Results with Llama-3. We expanded our evaluation beyond Qwen2.5 and conducted additional experiments using Llama-3.1-70B-Instruct and Llama-3.1-8B-Instruct pair (non-quantized version), with model weights distributed on 8 H20 GPUs. The results for both single-draft and multi-draft setups are shown in Table˜4(a).
Extended Results with EAGLE-3. To further demonstrate compatibility with advanced decoding systems, we integrated HSD into the SOTA EAGLE-3-LLaMa3.1-Instruct-8B (draft length $\gamma=7$ by default) by replacing its tokenwise verifier in Table 4(b). As mentioned in Section˜2, EAGLE-3 always accepts the first draft token to maximize throughput, and we follow the same setting for a fair comparison. Replacing the baseline verifier with HSD consistently improves decoding efficiency.
@@ E Extended Experiments
F Integration with Recursive Reject Sampling in the Multi-Draft Setup

### v2

# arXiv:2601.05724v2[cs.AI] 02 Mar 2026

@@ 1 Introduction
Inference speed has become paramount for Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Bai et al., 2023), which generate text auto-regressively. Recent advances in test-time scaling (OpenAI, 2024; Guo et al., 2025; Yu et al., 2025; Peng et al., 2025) have further underscored its importance. While techniques like pruning (Frankle and Carbin, 2018; Sun et al., 2023a) and quantization (Shen et al., 2020; Xiao et al., 2023) improve efficiency but sacrifice performance, Speculative Decoding (Leviathan et al., 2023) achieves speedups while preserving the target model’s distribution, making it a particularly appealing alternative. It adopts a smaller model to make proposals and a larger model to select from them with a grounded verification strategy. Most approaches prioritize the drafting phase, but further gains face diminishing returns. Driven by the verification bottleneck, recent methods (Cai et al., 2024; Zhou et al., 2024; Narasimhan et al., 2024) trade off fidelity for speed, relying on task-specific tuning; their performance typically remains constrained to carefully curated scenarios.
@@ 2 Related Work
Drafting Phase. Drafting methods can be grouped into three categories: (1) Single-draft. Early SD methods (Leviathan et al., 2023) inspired PaSS (Monea et al., 2023) and Draft&Verify (Zhang et al., 2024), improving efficiency via multi-token generation or selective layer skipping. GLIDE (Du et al., 2024) (shared KV-cache) offers further speedups but requires task-specific tuning. (2) Retrieval-based. LLM-A (Yang et al., 2023) and ReST (He et al., 2023) generate drafts from reference texts, potentially reducing latency, but face database limitations, distribution gaps, and reliance on greedy decoding. (3) Multi-draft. Tree-attention frameworks—SpecInfer (Miao et al., 2024), Medusa (Cai et al., 2024), and Eagle (Li et al., 2024; Fan et al., 2026)—expand many branches, quickly exhausting memory. Medusa and Eagle also predict drafts from the target model’s hidden features rather than a separate draft model, further boosting speed but requires task-specific tuning.
Verification Phase. Verification methods trade fidelity for speed. Lossless approaches (Sun et al., 2023b; Yang et al., 2024; Hu et al., 2025) guarantee exact recovery but are costly. Block Verification (Sun et al., 2024) partially alleviates this bottleneck but offers limited improvement and low interpretability and integrity. Lossy methods—including BiLD (Kim et al., 2023), MTAD (Qin et al., 2025), DistillSpec (Zhou et al., 2024), Medusa-2 (Cai et al., 2024), SpecCascade (Narasimhan et al., 2024) and CoS (Fu et al., 2025) increase speed but compromise distribution fidelity and require task-specific tuning. In addition, Medusa and EAGLE always accept the first draft token to improve throughput, trading off exact recovery of the target distribution.
@@ 5.3 Computational Efficiency
The verification stage in HSD adds negligible overhead compared to the savings from reduced target model forward passes. Thanks to parallelized computations across both the vocabulary and draft positions, HSD is nearly as efficient as tokenwise verification. Runtime measurements (Appendix H) show that verification accounts for less than 1% of total decoding time, with the majority still spent on draft and target forward passes. These results demonstrate that HSD is not only theoretically lossless but also practically efficient, as further confirmed by our experiments in Section˜6.
@@ 6.1 Experiment Setting
Baselines and Metrics. We compare two lossless verification methods—Token-wise and Block-wise—using two metrics: Block Efficiency (tokens/step) and Decoding Speed (tokens/second). Block Efficiency measures the average tokens generated per serial call to the target model, reflecting intrinsic efficiency independent of hardware. Decoding Speed indicates tokens produced per second for practical reference. Additional details and extended evaluations are in Section˜E.
@@ 6.2 Experiment Results
Main results.  Table 1 summarizes the performance of HSD across datasets and model scales using the Qwen2.5 suite (0.5B as draft,14B, 32B, and 72B as targets). Overall, HSD consistently improves both Block Efficiency (BE) and Decoding Speed (DS) relative to Tokenwise and Blockwise verification. For GSM8K, the gains are stable across scales, with BE improvements of 5.2%–5.4% at 14B/32B and 3.3% at 72B, accompanied by DS increases of up to 10.7%. On HumanEval, the effect is more pronounced: BE rises by 9.5% and 12.3% at 14B and 32B, while DS improves by 9.3% and 11.4%; even at 72B, HSD maintains positive margins (3.3% BE, 4.5% DS). For CNN/DailyMail, the improvements are moderate but consistent, with BE gains of 4.2%–8.4% and DS gains of 3.4%–7.2%. On average, HSD provides consistent advantages over Tokenwise and Blockwise verification, with improvements of approximately 6.2% in BE and 6.7% in DS.
Multi-draft. To demonstrate the compatibility of HSD, we compare it with token-wise verificaiton in a multi-draft setting. For simplicity—and without loss of generality—we adopt Recursive Reject Sampling (RRS) with replacement (Yang et al., 2024) as the baseline for its scalability and independence from complex tree attention mechanisms. Notably, since it is not straightforward to extend blockwise verification to the multi-draft setup, we omit it from our comparison. We evaluated multi-draft generation with 11 candidate drafts in Table 2, and HSD yields an average 5.9% improvement in Block Efficiency and 4.7% improvement in Decoding Speed over token-wise decoding.
@@
(a) Ablation on temperature ($\gamma=10$).
@@
(b) Ablation on draft lengths ($t=1$).
@@
Extended Results. We conducted additional experiments using Llama-3.1-70B-Instruct and Llama-3.1-8B-Instruct pair (non-quantized version), with model weights distributed on 8 H20 GPUs. The results are shown in Table˜4(a). Moreover, we integrated HSD into the SOTA EAGLE-3-LLaMa3.1-Instruct-8B ($\gamma=7$) by replacing its tokenwise verifier in Table 4(b). Following EAGLE-3, we accept at least the first draft token for a fair comparison. Note that EAGLE-3 utilizes top-K sampling for drafting, making all draft probabilities equal to 1. In this case, any verification method theoretically degenerates into the same behavior and the observed gain in block efficiency of HSD is likely influenced by sampling stochasticity and floating-point precision. However, the observed significant practical speedup in decoding speed is expected, since our implementation (see Section˜F) avoids the explicit loops in EAGLE’s implementation of tokenwise verification.
@@ E Extended Experiments
F Python Implementation
We provide the Python implementation of our Hierarchical Speculative Decoding (HSD) algorithm in Listing˜2, which builds upon the token-wise speculative decoding approach from Hugging Face Wolf et al. [2020] Transformers v4.46.3, shown in Listing˜1 for comparison. Following Hugging Face, our implementation eliminates the use of an explicit for-loop by leveraging an equivalent masking mechanism: we perform parallel sampling across all positions to determine whether to accept or reject subsequences of varying lengths, and then select the longest accepted prefix as the final output.
Listing 1 Tokenwise Speculative Decoding (SD) SD.py
1import torch
2
3def SD(candidate_input_ids, candidate_logits, new_logits):
4 """
5 Args:
6 candidate_input_ids (Tensor): Token IDs from the draft model. Shape: [batch_size, seq_len]
7 candidate_logits (Tensor): Logits from the draft model. Shape: [batch_size, seq_len, vocab_size]
8 new_logits (Tensor): Logits from the target model. Shape: [batch_size, seq_len, vocab_size]
9 Returns:
10 n_matches (int): Number of accepted tokens from the draft model.
11 valid_tokens (Tensor): Accepted token prefix with one new token sampled. Shape: [batch_size, n_matches+1]
12 """
13
14 # Convert logits to probabilities
15 q = candidate_logits.softmax(dim=-1)
16 p = new_logits.softmax(dim=-1)
17
18 candidate_length = candidate_logits.shape[1]
19 new_candidate_input_ids = candidate_input_ids[:, -candidate_length:]
20
21 # Extract token-wise probabilities for the candidate tokens
22 q_i = q[:, torch.arange(candidate_length), new_candidate_input_ids].squeeze(1)
23 p_i = p[:, torch.arange(candidate_length), new_candidate_input_ids].squeeze(1)
24
25 probability_ratio = p_i / q_i
26 is_accepted = torch.rand_like(probability_ratio) <= probability_ratio
27
28 # assuming batch size = 1
[…]
30
31 # Next token selection: if there is a rejection, adopt the resampling distribution.
32 if n_matches < candidate_length:
33 p_n_plus_1 = p[:, n_matches, :]
34 q_n_plus_1 = q[:, n_matches, :]
35 p_prime = torch.clamp((p_n_plus_1 - q_n_plus_1), min=0)
36 p_prime.div_(p_prime.sum())
37 else:
38 p_prime = p[:, n_matches, :]
39
40 # Ensure we don’t generate beyond max_len or an EOS token.
41 if is_done_candidate[0] and n_matches == candidate_length:
42
43 # Output length is assumed to be ‘n_matches + 1‘. Since we won’t generate another token with the target model
44 # due to acceptance on EOS we fix ‘n_matches‘
45 n_matches -= 1
46 valid_tokens = candidate_input_ids[:, -candidate_length:]
47
48 else:
49 # Next token selection: if there is a rejection, adjust the distribution from the main model before sampling.
50 # The selected tokens include the matches (if any) plus the next sampled tokens
51 if n_matches > 0:
52 if n_matches < candidate_length:
53 valid_tokens = candidate_input_ids[:, -candidate_length:n_matches - candidate_length]
54 if not stop(valid_tokens, scores=None):
55 t = torch.multinomial(p_prime, num_samples=1)
56 valid_tokens = torch.cat(
57 (valid_tokens, t), dim=-1)
58 else:
59 n_matches = n_matches-1
60 else:
61 valid_tokens = candidate_input_ids[:, -candidate_length:]
62 if not stop(valid_tokens, scores=None):
63 t = torch.multinomial(p_prime, num_samples=1)
64 valid_tokens = torch.cat(
65 (valid_tokens, t), dim=-1)
66 else:
67 n_matches = n_matches -1
68 else:
69 t = torch.multinomial(p_prime, num_samples=1)
70 valid_tokens = t
71
72 return valid_tokens, n_matches
Listing 2 Hierarchical Speculative Decoding (HSD) HSD.py
1import torch
2
3def HSD(candidate_input_ids, candidate_logits, new_logits):
4 """
5 Args:
6 candidate_input_ids (Tensor): Token IDs from the draft model. Shape: [batch_size, seq_len]
7 candidate_logits (Tensor): Logits from the draft model. Shape: [batch_size, seq_len, vocab_size]
8 new_logits (Tensor): Logits from the target model. Shape: [batch_size, seq_len, vocab_size]
9 Returns:
10 n_matches (int): Number of accepted tokens from the draft model.
11 valid_tokens (Tensor): Accepted token prefix with one new token sampled. Shape: [batch_size, n_matches+1]
12 """
13
14 # Convert logits to probabilities
15 q = candidate_logits.softmax(dim=-1)
16 p = new_logits.softmax(dim=-1)
17 candidate_length = candidate_logits.shape[1]
18 new_candidate_input_ids = candidate_input_ids[:, -candidate_length:]
19
20 # Extract token-wise probabilities for the candidate tokens
21 q_i = q[:, torch.arange(candidate_length), new_candidate_input_ids].squeeze(1)
22 p_i = p[:, torch.arange(candidate_length), new_candidate_input_ids].squeeze(1)
23
24 # Compute cumulative joint probabilities for draft and target model
25 q_prev = torch.roll(q_i, shifts=1, dims=1)
26 q_prev[:, 0] = 1.0
27 q_cumprod = torch.exp(torch.log(q_prev).cumsum(dim=1)).unsqueeze(-1)
28 q_next = q_cumprod * q[:, :candidate_length]
29 p_prev = torch.roll(p_i, shifts=1, dims=1)
30 p_prev[:, 0] = 1.0
31 p_cumprod = torch.exp(torch.log(p_prev).cumsum(dim=1)).unsqueeze(-1)
32
33 # Constrain p_cumprod with q_cumprod for computing the capped resampling distribution
34 ratio = p_cumprod / q_cumprod
35 previous_max = 1
36 new_p_previous = torch.ones_like(p_cumprod).to(p_cumprod.device)
37 for k in range(candidate_length):
38 if ratio[:, k] > previous_max:
39 previous_max = ratio[:, k]
40 new_p_previous[:, k] = p_cumprod[:, k] / previous_max
41 p_next = new_p_previous * p[:, :candidate_length]
42
43 # Construct resampling distribution p’
44 diffs = p_next - q_next
45 p_plus = torch.clamp(diffs, min=0.0)
46 p_minus = torch.clamp(-diffs, min=0.0)
47 p_primes = p_plus / torch.maximum(p_plus.sum(dim=-1, keepdim=True), p_minus.sum(dim=-1, keepdim=True))
48
49 # Step-back probability: reject prefix with 1 - mass of p’
50 step_back_probs = 1 - p_primes.sum(dim=-1)
51 step_back = torch.rand_like(step_back_probs) < step_back_probs
52
53 # Find first position to stop (from the end)
54 if step_back.all():
55 stop_positions = 0
56 else:
[…]
58
59 # Mask to decide which tokens are accepted
60 select = torch.zeros_like(step_back).to(step_back.device)
61
62 # apply cumprod on the ratio instead of the raw probabilities to avoid underflow
63 probability_ratio = (p_i / q_i).cumprod(1).unsqueeze(-1)
64 is_accepted = torch.rand_like(probability_ratio) <= probability_ratio
65
66 # only decide to accept or not at the last position based on the joint probability ratio
67 # assign 0 to all positions when the full draft is rejected, otherwise assign 1 to the rest of the positions
[…]
69 is_accepted = 1 - torch.cumsum(select, dim=-1)
70
71 #### assume batch_size=1 for the current implementation
72 n_matches = is_accepted.sum().item()
Listing 3 Hierarchical Speculative Decoding HSD.py (cont.)
1 if is_done_candidate[:] and n_matches == candidate_length:
2 # Output length is assumed to be ‘n_matches + 1‘. Since we won’t generate another token with the target model
3 # due to acceptance on EOS we fix ‘n_matches‘
4 n_matches -= 1
5 # valid_tokens = new_candidate_input_ids[:, : n_matches + 1]
6 valid_tokens = candidate_input_ids[:, -candidate_length:]
7
8 else:
9 # Next token selection: if there is a rejection, adjust the distribution from the main model before sampling.
10 gamma = candidate_length
11 p_n_plus_1 = p[:, candidate_length, :]
12 if n_matches < gamma:
13 p_prime = p_primes[:, n_matches]
14 p_prime = p_prime/p_prime.sum(-1, keepdim=True)
15 else:
16 p_prime = p_n_plus_1
17
18 # The selected tokens include the matches (if any) plus the next sampled tokens
19 # because if n_matches=0, we add one resampled token for sure, if n_matches=10, we add one more for sure
20 # as well, because the previous if checked not stop and n_matches-candidate_length will be 0 causing problem
21 if n_matches > 0 and n_matches<candidate_length:
22 valid_tokens = candidate_input_ids[:, -candidate_length:n_matches-candidate_length]
23 if not stop(candidate_input_ids[:, :n_matches-candidate_length], scores=None):
24 t = torch.multinomial(p_prime, num_samples=1)
25 valid_tokens = torch.cat(
26 (valid_tokens, t), dim=-1)
27 else:
28 n_matches = n_matches-1
29 else:
30 t = torch.multinomial(p_prime, num_samples=1)
31 if n_matches==0:
32 valid_tokens = t
33 else:
34 valid_tokens = candidate_input_ids[:, -candidate_length:]
35 valid_tokens = torch.cat(
36 (valid_tokens, t), dim=-1)
37
38 return valid_tokens, n_matches
@@ 9 Returns:
G Integration with Recursive Reject Sampling in the Multi-Draft Setup